OpenStudy (anonymous):

Use inductive reasoning to find the next term in the sequence 1, 2, 4, 7, 11, 16… What is the next term? (Points : 1) 18 19 22 26

OpenStudy (solomonzelman):

I will try to give you a little hint. (1st difference) \(\large\color{black}{ \displaystyle \color{red}{{\rm d}_1=}a_2-a_1=2-1=\color{red}{1} \\[0.3em] }\) (2nd difference) \(\large\color{black}{ \displaystyle \color{red}{{\rm d}_2=}a_3-a_2=4-2=\color{red}{2} \\[0.3em] }\) (3rd difference) \(\large\color{black}{ \displaystyle\color{red}{{\rm d}_3=}a_4-a_3=7-4=\color{red}{3} \\[0.3em] }\) (4th difference) \(\large\color{black}{ \displaystyle\color{red}{{\rm d}_4=}a_5-a_4=11-7=\color{red}{4} \\[0.3em] }\) (5th difference) \(\large\color{black}{ \displaystyle\color{red}{{\rm d}_5=}a_6-a_5=16-11=\color{red}{5} \\[0.3em] }\)

OpenStudy (solomonzelman):

see that the difference is growing by 1, and in fact, and nth difference is "n" (where n is integer of course, since n denotes not only the difference, but the order of the term)

OpenStudy (anonymous):

i dont understand?

OpenStudy (solomonzelman):

When you just say "I don't understand" that doesn't tell me what your concern is. Anything specific?

OpenStudy (anonymous):

like..i dont understand how i would find the next term?

OpenStudy (loser66):